# DEFAME: Dynamic Evidence-based FAct-checking with Multimodal Experts

### v2

@@ 4.1 Datasets
We evaluate DEFAME on three well-known fact-checking datasets, representative of three distinct areas in fact-checking literature, and one novel benchmark comprised of claims after October 2023.
@@
AVeriTeC (Schlichtkrull et al., 2024b) is a popular text-only real-world-based benchmark. The development set consists of $500$ claims: $305$ [Uncaptioned image]Refuted, $122$ [Uncaptioned image]Supported, $35$ [Uncaptioned image]NEI (Not Enough Information), and $38$ claims with the [Uncaptioned image]C/CP (Conflicting/Cherrypicking) label that designates claims with conflicting evidence or claims that are technically true but taken out of context. We retrieve evidence from the benchmark-complementary Knowledge Base (KB), which contains the necessary evidence along with approximately $1,000$ unrelated resources to simulate open web search. Thus, for AVeriTeC, the Web Search Tool does not utilize the Serper API but searches for documents via a semantic search. Each query to the KB is encoded using gte-base-en-v1.5 (Alibaba-NLP, 2024); the closest documents to the search query are retrieved via $k$-nearest neighbor. We evaluate DEFAME’s performance by computing accuracy over all $4$ classes.
Dataset GPT-4o GPT-4o mini LLaVA-1V
VERITE 84.5 67.1 59.3
MOCHEG 59.5 55.5 42.1
AVeriTeC 74.0 59.2 40.0
ClaimReview2024+ 67.0 53.7 36.9
Table 2: DEFAME accuracy for three different backbone models (GPT-4o, GPT-4o mini, LLaVA-1V) on the four benchmarks.
@@
MOCHEG (Yao et al., 2023) is a multimodal dataset. Its textual claims are paired with multimodal text-image evidence and sourced from PolitiFact [^7] https://www.politifact.com/ and Snopes [^8] https://www.snopes.com/. The test set contains $2,001$ unique claims, out of which $1,689$ contain a final ruling. The ruling is used to assess the quality of generated justifications. Thus, we choose this subset as our test set. It includes $667$ [Uncaptioned image]Refuted, $522$ [Uncaptioned image]NEI, and $500$ [Uncaptioned image]Supported claims. We evaluate our performance using the micro $F1$-score for comparability with prior work while also reporting various metrics to assess justification quality in Appendix I.
The VERITE (Papadopoulos et al., 2024b) benchmark is designed for image-text claim verification focusing on the Out-Of-Context (OOC) scenario. It features claims with textual and image components, with some samples originating from fact-checking platforms and others obtained by swapping either an image or parts of a caption. We use the entire VERITE dataset totaling $1,001$ samples for evaluation.99913131313 claims were excluded due to issues in VERITE’s construction process, reducing the dataset from $1014$ to $1001$ samples. The samples are categorized into $338$ true, $325$ OOC, and $338$ miscaptioned claims [^10] These differ in the way they were constructed, but both result in out-of-context images.. We follow Papadopoulos et al. (2024a) and report “[Uncaptioned image]True vs. [Uncaptioned image]Miscaptioned” and “[Uncaptioned image]True vs. [Uncaptioned image]OOC” accuracy. We further merge the two fabricated classes into one and report the “[Uncaptioned image]True vs. [Uncaptioned image]False” accuracy.
Finally, we construct ClaimReview2024+, a novel dataset with claims sourced from ClaimReview [^11] https://www.claimreviewproject.com/. This dataset aims to avoid the data leakage issue, i.e., MLLMs are likely pretrained on significant portions of claims and correlated fact-check articles from the previous three benchmarks. ClaimReview2024+ dataset consists of $300$ claims sourced after the October 2023 knowledge cutoff of GPT-4o. That way, we simulate a more realistic scenario where DEFAME is confronted with new events not encoded in the MLLM’s parameters. The dataset is subdivided into $160$ unimodal and $140$ multimodal claims, adhering to a four-class labeling scheme inspired by AVeriTeC: [Uncaptioned image]Supported, [Uncaptioned image]Refuted, [Uncaptioned image]Misleading, and [Uncaptioned image]NEI. This labeling also closely follows the classification by the fact-checking organizations that contribute to ClaimReview. Additional details about the dataset are included in Appendix J.
@@ 4.2 Model and Configuration
We chose GPT-4o and GPT-4o mini as the backbone of DEFAME since they are the current state-of-the-art MLLMs. To account for open-source MLLMs, we also test LLaVa-OneVision (1V) (Li et al., 2024a). DEFAME includes the MLLM without any fine-tuning, with temperature set to $0.01$ and top-$p$ to $0.9$ to control response diversity. We limit the number of images per scraped web page to a maximum of $32$ to avoid an excessive flood of images. DEFAME processes interleaved text-image inputs, preserving the original position of images within the text context, but any input exceeding the MLLM’s maximum context window of $128\,k$ tokens is truncated. Table 2 shows the performance of the different backbones.
@@ 4.3 Comparison to the State-of-the-Art
In Table 3, we present our evaluation results compared to State-of-the-Art (SOTA) methods and two baselines: GPT-4o, which directly generates a verdict (“Determine the claim’s veracity by picking one of the following decision options…”), and GPT-4o Chain-of-Thought (CoT), which also relies solely on parametric knowledge but is instructed to perform step-by-step reasoning before generating the verdict. DEFAME achieves an accuracy of $74.0\%$ on the AVeriTeC benchmark and surpasses the previous SOTA (Cao et al., 2023) that deploys a CoT LLM approach. With an overall accuracy of $84.5\%$ on VERITE, DEFAME ranks $32.7$ percentage points above prior best result (Papadopoulos et al., 2024a) in terms of “[Uncaptioned image]True vs. [Uncaptioned image]Miscaptioned” accuracy. It performs competitively on the “[Uncaptioned image]True vs. [Uncaptioned image]OOC” accuracy, leading to an overall accuracy gain of $26.5$ percentage points. On MOCHEG, our framework achieves a micro $F1$-score of $59.5\%$, replacing the MetaSumPerceiver (Chen et al., 2024) as the new SOTA fact-checking system. Importantly, no prior work has demonstrated the ability to simultaneously address such diverse tasks as we do here.
Model AVeriTeC MOCHEG VERITE CR+
Acc F1 T/OOC T/MC T/F Acc
CFR (2024) 60.0 - - - - -
GPT-CoT (2023) 65.6 - - - - -
LVLM4FV (2024) - 45.1 - - - -
MetaSum (2024) - 48.6 - - - -
CHASMA (2024b) - - 74.4* 59.3* 52.1 -
AITR (2024a) - - 82.7 51.8 58.0* -
GPT-4o 61.4 53.5 68.9 71.6 78.0 34.7
GPT-4o CoT 60.6 49.6 73.6 76.2 79.7 36.6
DEFAME 74.0 59.5 79.8 84.5 84.5 67.0
Table 3: Comparison of our method with prior works across datasets. Best scores are in bold, second best are underlined. Values marked with * had to be derived from the reported numbers. CR+ = ClaimReview2024+.
@@
The results on our new ClaimReview2024+ dataset challenge the notion that vanilla LLMs are reliable fact-checking systems. While performance on established benchmarks is strong, ClaimReview2024+ reveals a drastic drop for the two GPT-4o baselines, whereas DEFAME maintains accuracy comparable to its performance on AVeriTeC. This suggests that DEFAME’s evidence retrieval mitigates the temporal dependence of its backbone model. The confusion matrices (Figure 4) illustrate distinct error patterns across models on the ClaimReview2024+ dataset. Pure GPT-4o overpredicts [Uncaptioned image]NEI but outperforms its Chain-of-Thought counterpart in identifying [Uncaptioned image]Refuted claims. GPT-4o CoT also acknowledges the lack of evidence but shows a stronger tendency to commit to [Uncaptioned image]Misleading. In contrast, DEFAME exhibits a more balanced confusion matrix, with its primary challenge lying in differentiating [Uncaptioned image]Misleading and [Uncaptioned image]Refuted —a relatively uncritical ambiguity among the possible misclassifications. Its result on the [Uncaptioned image]Supported class is far superior to the two baselines.
@@
For example, when confronted with the claim, “Slovakian Prime Minister Robert Fico being dragged into a car after being shot,” (Fig. 3) the GPT-4o CoT baseline dismisses it as false due to a lack of evidence. In contrast, DEFAME conducts a reverse image search and web search, finds the image, including the corresponding articles on CNN, Vatican News, and Al Jazeera, and concludes that “the evidence confirms that Fico was indeed shot on May 15, 2024, in Handlová, Slovakia, and subsequently moved into a car after the incident.” Even though the geolocation suggested Poland and the Czech Republic, DEFAME correctly marginalized this finding and predicted the correct verdict.
Further analysis is provided in Appendix D.
(a) ClaimReview - GPT-4o
(b) ClaimReview - GPT-4o CoT
(c) ClaimReview - DEFAME
Figure 4: Confusion matrices for GPT-4o, GPT-4o CoT, and DEFAME on the ClaimReview2024+ dataset.
@@ 4.4 Ablation Study
MOCHEG VERITE
Model Variant F1 T/F (Acc.)
DEFAME 59.5 84.5
w/o Geolocation 58.3 80.6
w/o Reverse Search 58.2 73.7
w/o Image Search 57.8 81.4
w/o Web Search 42.0 81.8
Single Turn 47.7 82.8
Static Actions 58.7 83.0
w/o Develop 57.4 83.8
Unimodal Develop 56.1 82.0
Table 4: Ablation study results for different model variants on the MOCHEG and VERITE datasets. “True vs. False(= MC + OOC)” accuracy is reported for VERITE. Best scores are marked in bold, second best are underlined.
@@
Static Actions: This variant fixes all actions into a static action schedule. Each tool is executed exactly once, bypassing the dynamic planning stage.
W/o Develop Stage: Is intermediate reasoning useful? We test a variant that excludes the Develop Stage, jumping from the evidence retrieval immediately to the judgment.
Unimodal Develop Stage: Does it help to carry images through the stages and iterations, or is textual reasoning sufficient? This part-unimodal variant can retrieve evidence images but cannot pass them on to future stages.
The results of the ablation study are displayed in Table 4. Analyzing the contribution of each tool, we find that DEFAME benefits from all four tools. While Web Search and Image Search are important for both datasets, Geolocation and Reverse Image Search yield little gain on MOCHEG. This is likely caused by the benchmark’s focus on textual claims. On the other hand, the images included in the VERITE claims can be immediately scrutinized with Geolocation and Reverse Image Search tools, thereby increasing these tools’ relevance for text-image claims.
The single-turn variant of DEFAME lags behind the multi-turn version, especially on the MOCHEG benchmark. This suggests that for MOCHEG follow-up retrievals are important. Our findings also highlight the significant impact of the Develop Stage and the integration of visual evidence from tool outputs into the fact-check report. This multimodal enhancement yields a measurable performance boost, improving results by around 3 percentage points on MOCHEG and VERITE. Additionally, our experiments indicate that a static action routine leads to inferior accuracy as well as significantly higher computational costs (cf. Appendix B).
@@ 4.5 Adaptation to the AVeriTeC Challenge
Figure 5: Top 8 systems on the AVeriTeC Challenge test set, ranked by AVeriTeC score (in %).
The AVeriTeC dataset has recently run a challenge, where methods compete in terms of a new metric, AVeriTeC Score, aimed to capture accuracy and evidence correctness, see Schlichtkrull et al. (2024b) for details. To compute this metric, we adapt our approach to include question-answer (QA) pairs since the score measures the similarity between generated and gold QA pairs. See Appendix C for more details. Our adapted approach (coined DEFAME-QA) achieves stellar results in the AVeriTeC Challenge, shown in Table 5). Notably, both the original DEFAME and the adapted DEFAME-QA achieve the same accuracy on the development set.
@@ 4.6 Explainability Quality and Human Evaluation
Completeness: The verdict is sufficiently justified. That is, the included evidence allows one to derive the verdict.
Figure 6: Human assessment of Coherence and Completeness of DEFAME’s vs. GPT + CoT’s fact-checking reports
Participants were asked to rate the fact-checking reports w.r.t. the above criteria on a Likert scale from $1$ to $5$. In total, we collect $185$ ratings over $36$ claims sampled randomly but balanced from VERITE, MOCHEG, and AVeriTeC. Please refer to Appendix H for further details on the experiment. The results are shown in Fig. 6. The CoT baseline and DEFAME show no significant difference in output coherence, i.e., both produce mostly logical reports. This is expected, considering the current state of zero-shot LLMs in text generation. However, the two differ significantly in their ability to justify their verdicts. The results imply that DEFAME provides better justifications compared to bare MLLM prompting. This finding further challenges recent studies (Geng et al., 2024; Beigi et al., 2024; Shao et al., 2023; Li et al., 2024e) that suggest MLLMs can perform fact-checking without retrieving external evidence.

### v3

@@ 5.1 Datasets
Next to ClaimReview2024+, we evaluate DEFAME on three well-known fact-checking datasets, representative of distinct areas in fact-checking literature.
@@
AVeriTeC (Schlichtkrull et al., 2024b) is a popular text-only real-world-based benchmark. The development set consists of $500$ claims: $305$ [Uncaptioned image]Refuted, $122$ [Uncaptioned image]Supported, $35$ [Uncaptioned image]NEI (Not Enough Information), and $38$ claims with the [Uncaptioned image]C/CP (Conflicting/Cherrypicking) label that designates claims with conflicting evidence or claims that are technically true but lack context. We retrieve evidence from the benchmark-complementary Knowledge Base (KB), which contains the necessary evidence along with approximately $1,000$ unrelated resources to simulate open web search. Thus, for AVeriTeC, the Web Search Tool does not utilize the Serper API but rather a semantic search, yielding $5$ results for each search. Each query to the KB is encoded using gte-base-en-v1.5 (Alibaba-NLP, 2024); the closest documents to the search query are retrieved via $k$-nearest neighbor. We report the accuracy over all $4$ classes.
@@
MOCHEG (Yao et al., 2023) features textual claims paired with text-image evidence. Its multimodal nature qualifies it as a benchmark for DEFAME. Out of the $2,001$ unique claims in the test set, we choose the $1,689$ claims that have a final ruling, useful to assess the quality of generated justifications (Appendix M). That subset includes $667$ [Uncaptioned image]Refuted, $522$ [Uncaptioned image]NEI, and $500$ [Uncaptioned image]Supported claims. We evaluate our performance using accuracy—equivalent to micro-F1 (Appendix J).
VERITE (Papadopoulos et al., 2024b) is an image-text verification benchmark focused on Out-Of-Context (OOC) scenarios. After removing $13$ incomplete instances, VERITE comprises $1,001$ samples, sourced partly from fact-checking platforms and partly generated by swapping images or altering captions. The dataset includes $338$ [Uncaptioned image]True, $325$ [Uncaptioned image]OOC, and $338$ [Uncaptioned image]Miscaptioned claims (OOC and miscaptioned claims differ in construction but both involve out-of-context imagery). Following Papadopoulos et al. (2024a), we report accuracy for “[Uncaptioned image]True vs. [Uncaptioned image]OOC” and “[Uncaptioned image]True vs. [Uncaptioned image]Miscaptioned,” as well as a merged “[Uncaptioned image]True vs. [Uncaptioned image]False” setup.
DEFAME backbone
Dataset     GPT-4o     GPT-4o mini     LLaVA-1V     Llama 4
AVeriTeC 70.5 68.8 49.3 67.0
MOCHEG 59.5 55.5 42.1 55.0
VERITE 84.5 67.1 59.3 72.3
ClaimReview2024+ 69.7 47.7 32.6 48.8
Table 2: DEFAME accuracy with different backbone MLLMs (columns) on the four benchmarks (rows).
Method AVeriTeC MOCHEG VERITE CR+
Acc Acc T/OOC T/MC T/F Acc
CFR (Sriram et al., 2024) 60.0 - - - - -
DeBERTa (Cao et al., 2023) 65.6 - - - - -
LVLM4FV (Tahmasebi et al., 2024) - 45.1 - - - -
MetaSum (Chen et al., 2024) - 48.6 - - - -
CHASMA (Papadopoulos et al., 2024b) - - 74.4* 59.3* 52.1 -
AITR (Papadopoulos et al., 2024a) - - 82.7 51.8 58.0* -
GPT-4o 62.1 ± 0.4 53.7 ± 0.4 70.4 ± 1.4 72.0 ± 0.8 78.7 ± 0.8 35.2 ± 0.9
GPT-4o CoT 61.0 ± 0.4 49.7 ± 0.3 74.1 ± 0.9 76.5 ± 0.4 80.0 ± 0.6 31.4 ± 4.5
DEFAME (Ours) 70.5 ± 0.6 59.2 ± 0.4 78.4 ± 1.0 83.3 ± 1.1 83.9 ± 0.5 69.7 ± 2.5
Table 3: Comparison of our method with prior state-of-the-art methods and baselines across datasets. Best scores are in bold, second-best are underlined. For DEFAME and the GPT-4o baselines, we report the mean over three runs along with the standard deviation. Values marked with * were derived from reported numbers. Blank cells indicate that the respective method could not be evaluated on the corresponding benchmark—either due to unavailability of publicly released code or incompatibility with the benchmark setup. CR+ = ClaimReview2024+.
@@ 5.2 Model and Configuration
We chose GPT-4o and GPT-4o mini as the backbone of DEFAME since they are the current state-of-the-art MLLMs. To account for open-source MLLMs, we also test LLaVa-OneVision (1V) (7B) (Li et al., 2024a), and Llama 4 Scout (Meta AI, 2025). DEFAME includes the MLLM without any fine-tuning, with temperature set to $0.01$ and top-$p$ to $0.9$ to control response diversity. We limit the number of images per scraped web page to a maximum of $32$ to avoid an excessive flood of images. DEFAME processes interleaved text-image inputs, preserving the original position of images within the text context, but any input exceeding the MLLM’s maximum context window is truncated accordingly. Table 2 shows the performance of the different backbones. LLaVa-OneVision (7B) struggled with formatting outputs consistently for downstream use, leading to a substantial performance drop. While Llama 4 Scout addressed this issue and achieved results comparable to GPT-4o mini—suggesting that open-source models are gradually closing the gap—GPT-4o still outperformed all other backbones by a wide margin, particularly on ClaimReview2024+ and VERITE.
@@ 5.3 Comparison to the State-of-the-Art
In Table 3, we present our evaluation results compared to State-of-the-Art (SOTA) methods and two baselines: GPT-4o, which directly generates a verdict (“Determine the claim’s veracity by picking one of the following decision options…”), and GPT-4o Chain-of-Thought (CoT), which also relies solely on parametric knowledge but is instructed to perform step-by-step reasoning before generating the verdict. For robust comparison, we run both baselines and our main variant three times on each dataset and report the mean performance. DEFAME achieves an accuracy of $70.5\%$ on the AVeriTeC benchmark and surpasses the previous SOTA (Cao et al., 2023) that deploys a CoT LLM approach. With an overall “[Uncaptioned image]True vs. [Uncaptioned image]False” accuracy of $83.9\%$ on VERITE, DEFAME ranks $25.9$ percentage points above prior best result (Papadopoulos et al., 2024a), similar for “[Uncaptioned image]True vs. [Uncaptioned image]Miscaptioned”. It performs competitively on the “[Uncaptioned image]True vs. [Uncaptioned image]OOC” accuracy. On MOCHEG, our framework achieves an average accuracy of $59.2\%$, replacing the MetaSumPerceiver (Chen et al., 2024) as the new SOTA fact-checking system.
Importantly, no prior work has demonstrated the ability to simultaneously address such diverse tasks as we do here—making DEFAME the most general AFC method as of now. Unlike prior methods that are limited to specific subtasks or modalities—such as OOC detection, text-only verification, or reliance on gold evidence—DEFAME performs robustly across all benchmarks without task-specific tuning or training data. This benchmark-spanning performance is made possible by DEFAME’s zero-shot, retrieval-based design, which does not depend on gold-labeled evidence, modality constraints, or dataset-specific preprocessing.
(a) GPT-4o
(b) GPT-4o CoT
(c) DEFAME
Figure 4: Confusion matrices on the ClaimReview2024+ dataset for GPT-4o, GPT-4o CoT, and DEFAME.
@@
The results on our new ClaimReview2024+ dataset challenge the notion that vanilla LLMs are reliable fact-checking systems. While performance on established benchmarks is strong, ClaimReview2024+ reveals a drastic drop for the two GPT-4o baselines, whereas DEFAME maintains accuracy comparable to its performance on AVeriTeC. This suggests that DEFAME’s evidence retrieval mitigates the temporal dependence of its backbone model. The confusion matrices (Figure 4) illustrate distinct error patterns across models on the ClaimReview2024+ dataset. Pure GPT-4o overpredicts [Uncaptioned image]NEI but outperforms its Chain-of-Thought counterpart in identifying [Uncaptioned image]Refuted claims. GPT-4o CoT also acknowledges the lack of evidence but shows a stronger tendency to commit to [Uncaptioned image]Misleading. In contrast, DEFAME exhibits a more balanced confusion matrix, with its primary challenge lying in differentiating [Uncaptioned image]Misleading and [Uncaptioned image]Refuted—a relatively uncritical ambiguity among the possible misclassifications. Its result on the [Uncaptioned image]Supported class is far superior to the two baselines. Refer to Appendix G for further confusion analysis.
Figure 5 exemplarily shows the claim “Slovakian Prime Minister Robert Fico being dragged into a car after being shot.” The GPT-4o CoT baseline refutes it due to missing evidence. In contrast, DEFAME retrieves supporting articles from CNN, Vatican News, and Al Jazeera via reverse image and web search, confirming that Fico was indeed shot on May 15, 2024, in Handlová, Slovakia, and subsequently moved into a car. Although the geolocation suggested Poland or the Czech Republic, DEFAME correctly marginalized this noise and predicted the correct verdict.
@@ 5.4 Ablation Study
MOCHEG VERITE CR+
Model Variant Acc. T/F (Acc.) Acc.
DEFAME 59.2 83.9 69.7
w/o Geolocation 58.3 80.6 65.7
w/o Reverse Search 58.2 73.7 64.0
w/o Image Search 57.8 81.4 63.7
w/o Web Search 42.0 81.8 59.7
Single Turn 47.7 82.8 63.3
w/o Planning 58.7 83.0 68.0
w/o Develop 57.4 83.8 67.0
Unimodal Develop 56.1 82.0 65.7
Table 4: Ablation study results for different model variants on the MOCHEG, VERITE and ClaimReview2024+ datasets. “[Uncaptioned image]True vs. [Uncaptioned image]False” accuracy is reported for VERITE. Best scores are marked in bold, second best are underlined.
@@
  • W/o Planning: This variant fixes all actions into a static action schedule. Each tool is executed exactly once, bypassing the dynamic planning stage.
  • W/o Develop Stage: Is intermediate reasoning useful? This variant excludes the Develop Stage, jumping from evidence retrieval immediately to judgment.
  • Unimodal Develop Stage: Is textual reasoning sufficient? This part-unimodal variant can retrieve evidence images but cannot pass them on to future stages.
Table 4 presents the results of our ablation study. We observe that DEFAME benefits from all four tools—with varying contributions: Web Search is critical for MOCHEG and ClaimReview2024+, likely because both contain only or many text-only claims, respectively. Geolocation and Reverse Image Search prove more useful for VERITE, where images are central and can be directly examined.
The single-turn variant of DEFAME performs notably worse than the multi-turn version, especially on MOCHEG, underscoring the importance of follow-up retrievals. The Develop Stage and the integration of visual evidence from tool outputs into the report yield a clear performance gain. Finally, removing the Planning Stage leads not only to reduced accuracy but also to significantly higher computational costs (see Appendix C).
@@ 5.5 Adaptation to the AVeriTeC Challenge
Figure 6: Top 8 (of 21) systems on the AVeriTeC Challenge test set, ranked by AVeriTeC score.
The AVeriTeC authors recently hosted a shared task (Schlichtkrull et al., 2024a), where systems competed based on the AVeriTeC Score (Schlichtkrull et al., 2024b), a metric that evaluates both claim accuracy and evidence correctness. To participate, we adapted our framework to produce question-answer (QA) pairs, as the metric relies on the similarity between generated and reference QA pairs. Our adapted method, submitted under the name InFact, achieved best performance in the AVeriTeC Challenge, as shown in Figure 6 (see also Appendix E and Rothermel et al. (2024) for details). Notably, both the original DEFAME and the adapted InFact variant achieve comparable accuracy on the development set, demonstrating that our system is not only robust across evaluation protocols but also flexible enough to support output formats required by the downstream tasks.
@@ 5.6 Explainability Quality and Human Evaluation
  • Completeness: The verdict is sufficiently justified: he included evidence allows one to derive the verdict.
Figure 7: Human assessment of Coherence and Completeness of DEFAME’s vs. GPT-4o CoT’s fact-checking reports on $36$ claims sampled from three benchmarks.
Participants were asked to rate the fact-checking reports w.r.t. the above criteria on a Likert scale from $1$ to $5$. In total, we collect $185$ ratings over $36$ claims sampled randomly but balanced from VERITE, MOCHEG, and AVeriTeC. Please refer to Appendix L for further details on the experiment. The results are shown in Fig. 7. The CoT baseline and DEFAME show no significant difference in output coherence, i.e., both produce mostly logical reports. This is expected, considering the current state of zero-shot LLMs in text generation. However, the two differ significantly in their ability to justify their verdicts. The results imply that DEFAME provides better justifications compared to bare MLLM prompting. This finding further challenges recent studies (Geng et al., 2024; Beigi et al., 2024; Shao et al., 2023; Li et al., 2024e) that suggest MLLMs can perform fact-checking without retrieving external evidence.